# When Deep Learning Meets Polyhedral Theory: A Survey

### v3

@@ 1.1 What neural networks can model
where the weights ${\bm{w}}^{l}_{i}\in\mathbb{R}^{n_{l-1}}$ and the bias $b^{l}_{i}\in\mathbb{R}$ are parameters of the unit. Those parameters can be aggregated across the layer as the matrix ${\bm{W}}^{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and the vector ${\bm{b}}^{l}\in\mathbb{R}^{n_{l}}$. The vector ${\bm{h}}^{l-1}\in\mathbb{R}^{n_{l-1}}$ represents the aggregated outputs from layer $(l-1)$. The activation function $\sigma^{l}:\mathbb{R}\rightarrow\mathbb{R}$ is applied by the units in layer $l$. These definitions implicitly assume that $n_{0}$ is the size of the network input ${\bm{x}}\in\mathbb{R}^{n_{0}}$ and that ${\bm{h}}^{0}$ and ${\bm{x}}$ are the same. Figure 2 illustrates the operation of a feedforward network as described above.
@@ 1.5 Scope of this survey and related work
For a historical perspective on neural networks, we recommend Schmidhuber (2015). For a recent and broad introduction to the fundamentals of deep learning, we recommend Zhang et al. (2023). For other forms of measuring model complexity in neural networks, we refer to Hu et al. (2021).
@@ 3 Piecewise Linear Structure and the Linear Regions of a Neural Network
Linear regions are fundamental when analyzing the behavior of a neural network. If we restrict the domain of a neural network to a linear region ${\mathbb{I}}\subseteq\mathbb{R}^{n_{0}}$, then the neural network behaves as an affine transformation ${\bm{y}}_{\mathbb{I}}:{\mathbb{I}}\rightarrow\mathbb{R}^{n_{L}}$ of the form ${\bm{y}}_{\mathbb{I}}({\bm{x}})={\bm{T}}{\bm{x}}+{\bm{t}}$ with a matrix ${\bm{T}}\in\mathbb{R}^{n_{L}\times n_{0}}$ and a vector ${\bm{t}}\in\mathbb{R}^{n_{L}}$ that are directly defined by the network parameters and the set of neurons that are activated by any input ${\bm{x}}\in{\mathbb{I}}$. For a small perturbation $\varepsilon$ to some input $\overline{{\bm{x}}}\in{\mathbb{I}}$ such that $\overline{{\bm{x}}}+\varepsilon\in{\mathbb{I}}$, the network output for $\bar{x}+\varepsilon$ is given by ${\bm{y}}_{\mathbb{I}}(\overline{{\bm{x}}}+\varepsilon)$. While it is possible that two adjacent regions defined in such way correspond to the same affine transformation, thinking of each linear region as having a distinct signature of active units makes it easier to analyze them.
@@ 3.1 The combinatorial aspect of linear regions
Consider a neural network with input $x$ from the domain $[0,1]$ and layers having 4 neurons with ReLU activation. With the choice of parameters described in the next paragraph, we obtain a zigzagging function with 4 slopes in the $[0,1]$ domain (see $F(x)$ in Figure 3). Each of the slopes of $F(x)$ defines a bijection between segments of the input —namely, $[0,0.25]$, $[0.25,0.5]$, $[0.5,0.9]$, and $[0.9,1.0]$— and the image $[0,1]$. The effect of repeating such structure in the second layer is that of composing the zigzagging function with itself (see $F(F(x))$ in Figure 3), with 4 slopes being produced within each of those 4 initial segments. Hence, the number of slopes —and therefore of linear regions— in the output of such a neural network with $L$ layers of activation functions is $4^{L}$, which implies an exponential growth on depth.
@@
Figure 3: Output functions $\{f_{i}(x)\}_{i=1}^{4}$ of the units in the first layer and combined outputs of the first two layers —$F(x)=f_{1}(x)-f_{2}(x)+f_{3}(x)-f_{4}(x)$ for the first and $F(F(x))$ for the second— of a neural network in which the number of linear regions grows exponentially on the depth, as described in Example 1.
@@
The second observation —that the number of linear regions can grow exponentially in the number of neurons alone— comes from the interplay between the parts of the input space in which each of the units are active, especially in higher-dimensional spaces. This is based on some geometric observations that we discuss in Section 3.3. Even for a shallow network —i.e., the number of layers being $L=1$— such a number of linear regions may approach $2^{n}$, which corresponds to every possible activation set ${\mathbb{S}}\subseteq\{1,\ldots,n\}$ defining a nonempty linear region. However, as we discuss later, that is not always the case due to architectural choices such as the number of layers and their width.
@@ 3.2 The algebra of linear regions
Such partitioning entails an overlap between adjacent linear regions when ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}+b_{i}^{l}=0$, i.e., at the boundary in which unit $i$ in layer $l$ is active in one region and inactive in another. Nevertheless, for any input $\overline{{\bm{x}}}$ associated with a point at such a boundary between two linear regions ${\mathbb{I}}_{1}$ and ${\mathbb{I}}_{2}$, it holds that ${\bm{y}}_{{\mathbb{I}}_{1}}(\overline{{\bm{x}}})={\bm{y}}_{{\mathbb{I}}_{2}}(\overline{{\bm{x}}})$ even if those affine transformations are not entirely identical since the output of the neural network is continuous. More importantly, such overlap implies that each term of $\mathcal{D}$ is defined using only equalities and nonstrict inequalities. In other words, each linear region corresponds to a proper closed polyhedra in the extended $({\bm{x}},{\bm{h}}^{1},\ldots,{\bm{h}}^{L})$ space due to the overlap of boundaries between pairs of adjacent linear regions. Consequently, those linear regions also define polyhedra if projected to the ${\bm{x}}$ space, since by using Fourier-Motzkin elimination (Fourier, 1826, Motzkin, 1936) we obtain a polyhedral description of the linear region in ${\bm{x}}$. If one of those polyhedra in the ${\bm{x}}$ space does not have an interior, which means that it is not full-dimensional in ${\bm{x}}$, then that polyhedron lies entirely on the boundary of other linear regions. In such a case, we do not regard it as a proper linear region. The interior of every full-dimensional linear region is disjoint.
@@ 3.3 The geometry of linear regions
Consider a neural network with domain ${\bm{x}}\in\mathbb{R}^{2}$ and a single layer having 3 neurons $\alpha$, $\beta$, and $\gamma$ with outputs given as follows: $h^{1}_{\alpha}=\max\{2.3x_{1}-1.9x_{2}+0.6,0\}$, $h^{1}_{\beta}=\max\{-0.9x_{1}-0.7x_{2}+4.8,0\}$, and $h^{1}_{\gamma}=\max\{0x_{1}+3x_{2}-5,0\}$. These neurons define the activation hyperplanes ($\alpha$) $02.3x_{1}-1.9x_{2}+0.6=$ 0, ($\beta$) $-0.9x_{1}-0.7x_{2}+4.8=0$, and ($\gamma$) $00x_{1}+3x_{2}-5=$ 0 in the ${\bm{x}}$ space, which are illustrated along with the activation sets of the linear regions in Figure 5.
@@
Consider a neural network with domain ${\bm{x}}\in\mathbb{R}^{2}$ and 2 layers having 2 neurons each —say neurons $\alpha$ and $\beta$ in layer 1, and neurons $\gamma$ and $\delta$ in layer 2— with outputs given as follows: $h^{1}_{\alpha}=\max\{2.3x_{1}-1.9x_{2}+1.5,0\}$, $h^{1}_{\beta}=\max\{-0.9x_{1}-0.7x_{2}+5,0\}$, $h^{2}_{\gamma}=\max\{0.4h^{1}_{1}-3.1h^{1}_{2}+4,0\}$, $h^{2}_{\delta}=\max\{-0.6h^{1}_{1}-1.6h^{1}_{2}+5,0\}$. These neurons define the activation hyperplanes ($\alpha$) $02.3x_{1}-1.9x_{2}+1.5=$ 0 and ($\beta$) $-0.9x_{1}-0.7x_{2}+5=0$ in the ${\bm{x}}$ space and the activation hyperplanes ($\gamma$) $00.4h^{1}_{1}-3.1h^{1}_{2}+4=$ 0 and ($\delta$) $-0.6h^{1}_{1}-1.6h^{1}_{2}+5=0$ in the ${\bm{h}}^{1}$ space, which are illustrated along with the activation sets of the linear regions in the first two plots of Figure 6. The third plot illustrates the linear regions jointly defined by the two layers in terms of the input space, or ${\bm{x}}$ space.
@@
By comparing side by side the first and the third plots of Figure 6, we can see how every linear region of a given layer of a neural network may be partitioned differently by the following layer. When defined in terms of the input space, or ${\bm{x}}$ space, the hyperplanes associated with the second layer differ across the linear regions defined by the first layer because each of those linear regions is associated with a different affine transformation from ${\bm{x}}$ to ${\bm{h}}^{1}$. Hence, the activation hyperplanes of layer $l$ may break each linear region from layer $l-1$ differently. To every linear region defined by the hyperplane arrangement in the ${\bm{h}}^{l-1}$ space there is a linear transformation ${\bm{h}}^{l-1}=\Omega^{{\mathbb{S}}^{l-1}}({\bm{W}}^{l-1}{\bm{h}}^{l-2}+{\bm{b}}^{l-1})$ to the points of that linear region based on the set of active neurons ${\mathbb{S}}^{l-1}$. Consequently, inputs in the ${\bm{h}}^{l-1}$ space that are associated with different linear regions are transformed differently to the ${\bm{h}}^{l}$ space, and therefore the form in which those linear regions are further partitioned by layer $l$ is not the same when seen from the ${\bm{h}}^{l-1}$ space.
Second, some combinations of activation sets of multiple layers do not correspond to linear regions even if the activation hyperplanes are in general position with respect to each layer. For each layer, the first two plots of Figure 6 show that every activation set corresponds to a nonempty region of the layer input. However, not every pair of such activation sets would define a nonempty linear region for the neural network. For example, the linear region of the first layer associated with the activation set ${\mathbb{S}}^{1}=\{\}$ defines a linear region in ${\bm{x}}$ which is always mapped to ${\bm{h}}^{1}=0$, and thus only corresponds to activation set ${\mathbb{S}}^{2}=\{\gamma,\delta\}$ in the second layer because both units are active for such input. Thus, no linear region in ${\bm{x}}$ is associated with only the units in sets {},$\{\},\{\gamma\}$, and $\{\delta\}$ being active —i.e., there is no linear region such that ${\mathbb{S}}^{1}\cup{\mathbb{S}}^{2}=\{\},\{\gamma\},\text{or}\{\delta\}$.
More generally, the number of units that are active on each linear region defined by the first layer also imposes a geometric limit to how that linear region can be further partitioned. If only one unit is active at a layer, that means that the output of the layer within that linear region has dimension 1, and, consequently, the subsequent hyperplane arrangements within that linear region are limited to a 1-dimensional space. For the network in the example, we thus expect no more than $\sum_{i=0}^{1}\binom{2}{i}=3$ linear regions being defined instead of $42^{2}=$ 4 when only one unit is active. In fact, that is precisely the number of subdivisions by the second layer of the linear region defined by activation set ${\mathbb{S}}^{1}=\{\beta\}$ from the first layer.
@@ 3.4 Piecewise linear representability
A natural follow-up question is whether one can do better than this logarithmic bound. Hertrich et al. (2023) conjecture that this is not true.
@@
To narrow down the path into proving this conjecture, Hertrich et al. (2023) show that this conjecture is equivalent to the following conjecture.
@@
At the time of writing this survey, this remains a conjecture. However, considerable progress has been made. Hertrich et al. (2023) proved the following first step in this direction.
@@
Later on, the following notable step was given for the case of networks with integer weights.
@@
Consider $f$ a positively homogeneous, convex, piecewise linear function (such as the maximum function). Such a function can be written as $,f(x)=\max\{a_{1}^{\top}x,$$\ldots,a_{p}^{\top}x\}$ for some family of vectors $\{a_{i}\}_{i=1}^{p}$; the Newton polytope $P_{f}$ is defined as $P_{f}:=\mathrm{conv}\{a_{1},\ldots,a_{p}\}$. Note that if $f(x)=\max\{0,x_{1},\ldots,x_{n}\}$ its Newton polytope is the unit simplex, i.e., $P_{f}=\Delta_{0}^{n}:=\mathrm{conv}\{0,e_{1},\ldots,e_{n}\}$ where $e_{i}$ is the $n$-th dimensional $i$-th unit vector.
Haase et al. (2023) leverage a strong connection between Newton polytopes and ReLU networks: that the $+$ and $\max$ operations over functions (which are the building blocks in ReLU networks) translate well to operations over polytopes, namely, $P_{f+g}=P_{f}+P_{g}$ and $P_{\max\{f,g\}}=\mathrm{conv}(P_{f}\cup P_{g})$. Then, to connect the functions representable by networks with $k$ hidden layers and families of polytopes, the authors define the following recursive family.
$\mathcal{P}^{\prime}_{k+1}$ :=$:=\{\mathrm{conv}(Q_{1}\cup Q_{2})\,\mid\,Q_{1},Q_{2}\in\mathcal{P}_{k}\}$
$\mathcal{P}_{k+1}$ :=$:=\{Q_{1}+\cdots+Q_{\ell}\,\mid\,Q_{1},\ldots,Q_{\ell}\in\mathcal{P}^{\prime}_{k+1}\}.$
@@
A positively homogeneous continuous piecewise linear function $f$ can be represented by an integral $k$ hidden-layer neural network if and only if $f=g-h$ for two convex, positively homogeneous continuous piecewise linear functions $g$ and $h$ with $P_{g},P_{h}\in\mathcal{P}_{k}$.

### v4

@@ 1.1 What neural networks can model
where the weights ${\bm{w}}^{l}_{i}\in\mathbb{R}^{n_{l-1}}$ and the bias $b^{l}_{i}\in\mathbb{R}$ are parameters of the unit. Those parameters can be aggregated across the layer as the matrix ${\bm{W}}^{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and the vector ${\bm{b}}^{l}\in\mathbb{R}^{n_{l}}$. The vector ${\bm{h}}^{l-1}\in\mathbb{R}^{n_{l-1}}$ represents the aggregated outputs from layer (l1)(l-1). The activation function $\sigma^{l}:\mathbb{R}\rightarrow\mathbb{R}$ is applied by the units in layer $l$. These definitions implicitly assume that $n_{0}$ is the size of the network input ${\bm{x}}\in\mathbb{R}^{n_{0}}$ and that ${\bm{h}}^{0}$ and $\bm{x}$} are the same. Figure 2 illustrates the operation of a feedforward network as described above.
@@ 1.5 Scope of this survey and related work
For a historical perspective on neural networks, we recommend Schmidhuber (2015). For a recent and broad introduction to the fundamentals of deep learning, we recommend Zhang et al. (2023a). For other forms of measuring model complexity in neural networks, we refer to Hu et al. (2021).
@@ 3 Piecewise Linear Structure and the Linear Regions of a Neural Network
Linear regions are fundamental when analyzing the behavior of a neural network. If we restrict the domain of a neural network to a linear region ${\mathbb{I}}\subseteq\mathbb{R}^{n_{0}}$, then the neural network behaves as an affine transformation ${\bm{y}}_{\mathbb{I}}:{\mathbb{I}}\rightarrow\mathbb{R}^{n_{L}}$ of the form ${\bm{y}}_{\mathbb{I}}({\bm{x}})={\bm{T}}{\bm{x}}+{\bm{t}}$ with a matrix ${\bm{T}}\in\mathbb{R}^{n_{L}\times n_{0}}$ and a vector ${\bm{t}}\in\mathbb{R}^{n_{L}}$ that are directly defined by the network parameters and the set of neurons that are activated by any input ${\bm{x}}\in{\mathbb{I}}$. For a small perturbation $\varepsilon$ to some input $\overline{{\bm{x}}}\in{\mathbb{I}}$ such that $\overline{{\bm{x}}}+\varepsilon\in{\mathbb{I}}$, the network output for $\bar{x}+\varepsilon$ is given by ${\bm{y}}_{\mathbb{I}}(\overline{{\bm{x}}}+\varepsilon)$. While it is possible that two adjacent regions defined in such way correspond to the same affine transformation, thinking of each linear region as having a distinct signature of active units makes it easier to analyze them. Lee et al. (2019) and Hanin and Rolnick (2019b) have both discussed such distinction, and Stargalla et al. (2025) present additional definitions for the term “linear region” found in the literature.
@@ 3.1 The combinatorial aspect of linear regions
Consider a neural network with input $x$ from the domain $[0,1]$ and layers having 4 neurons with ReLU activation. With the choice of parameters described in the next paragraph, we obtain a zigzagging function with 4 slopes in the $[0,1]$ domain (see $F(x)$ in Figure 3). Each of the slopes of $F(x)$ defines a bijection between segments of the input —namely, $[0,0.25]$, $[0.25,0.5]$, $[0.5,0.9]$, and [0.9,1.0][0.9,1.0]— and the image $[0,1]$. The effect of repeating such structure in the second layer is that of composing the zigzagging function with itself (see $F(F(x))$ in Figure 3), with 4 slopes being produced within each of those 4 initial segments. Hence, the number of slopes —and therefore of linear regions— in the output of such a neural network with $L$ layers of activation functions is $4^{L}$, which implies an exponential growth on depth.
@@
Figure 3: Output functions $\{f_{i}(x)\}_{i=1}^{4}$ of the units in the first layer and combined outputs of the first two layers $F(x)=f_{1}(x)-f_{2}(x)+f_{3}(x)-f_{4}(x)$ for the first and $F(F(x))$ for the second— of a neural network in which the number of linear regions grows exponentially on the depth, as described in Example 1.
@@
The second observation —that the number of linear regions can grow exponentially in the number of neurons alone— comes from the interplay between the parts of the input space in which each of the units are active, especially in higher-dimensional spaces. This is based on some geometric observations that we discuss in Section 3.3. Even for a shallow network —i.e., the number of layers being L=1L=1— such a number of linear regions may approach $2^{n}$, which corresponds to every possible activation set ${\mathbb{S}}\subseteq\{1,\ldots,n\}$ defining a nonempty linear region. However, as we discuss later, that is not always the case due to architectural choices such as the number of layers and their width.
@@ 3.2 The algebra of linear regions
Such partitioning entails an overlap between adjacent linear regions when ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}+b_{i}^{l}=0$, i.e., at the boundary in which unit $i$ in layer $l$ is active in one region and inactive in another. Nevertheless, for any input $\overline{{\bm{x}}}$ associated with a point at such a boundary between two linear regions ${\mathbb{I}}_{1}$ and ${\mathbb{I}}_{2}$, it holds that ${\bm{y}}_{{\mathbb{I}}_{1}}(\overline{{\bm{x}}})={\bm{y}}_{{\mathbb{I}}_{2}}(\overline{{\bm{x}}})$ even if those affine transformations are not entirely identical since the output of the neural network is continuous. More importantly, such overlap implies that each term of $\mathcal{D}$ is defined using only equalities and nonstrict inequalities. In other words, each linear region corresponds to a proper closed polyhedra in the extended $({\bm{x}},{\bm{h}}^{1},\ldots,{\bm{h}}^{L})$ space due to the overlap of boundaries between pairs of adjacent linear regions. Consequently, those linear regions also define polyhedra if projected to the $\bm{x}$} space, since by using Fourier-Motzkin elimination (Fourier, 1826, Motzkin, 1936) we obtain a polyhedral description of the linear region in $\bm{x}$}. If one of those polyhedra in the $\bm{x}$} space does not have an interior, which means that it is not full-dimensional in $\bm{x}$}, then that polyhedron lies entirely on the boundary of other linear regions. In such a case, we do not regard it as a proper linear region. The interiors of all full-dimensional linear regions are disjoint.
@@ 3.3 The geometry of linear regions
Consider a neural network with domain ${\bm{x}}\in\mathbb{R}^{2}$ and a single layer having 3 neurons $\alpha$, $\beta$, and $\gamma$ with outputs given as follows: $h^{1}_{\alpha}=\max\{2.3x_{1}-1.9x_{2}+0.6,0\}$, $h^{1}_{\beta}=\max\{-0.9x_{1}-0.7x_{2}+4.8,0\}$, and $h^{1}_{\gamma}=\max\{0x_{1}+3x_{2}-5,0\}$. These neurons define the activation hyperplanes ($\alpha$) $2.3x_{1}-1.9x_{2}+0.6=0$, ($\beta$) $0.9x_{1}-0.7x_{2}+4.8=0$, and ($\gamma$) $0x_{1}+3x_{2}-5=0$ in the $\bm{x}$} space, which are illustrated along with the activation sets of the linear regions in Figure 5.
@@
Consider a neural network with domain ${\bm{x}}\in\mathbb{R}^{2}$ and 2 layers having 2 neurons each —say neurons $\alpha$ and $\beta$ in layer 1, and neurons $\gamma$ and $\delta$ in layer 2— with outputs given as follows: $h^{1}_{\alpha}=\max\{2.3x_{1}-1.9x_{2}+1.5,0\}$, $h^{1}_{\beta}=\max\{-0.9x_{1}-0.7x_{2}+5,0\}$, $h^{2}_{\gamma}=\max\{0.4h^{1}_{1}-3.1h^{1}_{2}+4,0\}$, $h^{2}_{\delta}=\max\{-0.6h^{1}_{1}-1.6h^{1}_{2}+5,0\}$. These neurons define the activation hyperplanes ($\alpha$) $2.3x_{1}-1.9x_{2}+1.5=0$ and ($\beta$) $0.9x_{1}-0.7x_{2}+5=0$ in the $\bm{x}$} space and the activation hyperplanes ($\gamma$) $0.4h^{1}_{1}-3.1h^{1}_{2}+4=0$ and ($\delta$) $0.6h^{1}_{1}-1.6h^{1}_{2}+5=0$ in the ${\bm{h}}^{1}$ space, which are illustrated along with the activation sets of the linear regions in the first two plots of Figure 6. The third plot illustrates the linear regions jointly defined by the two layers in terms of the input space, or $\bm{x}$} space.
@@
By comparing side by side the first and the third plots of Figure 6, we can see how every linear region of a given layer of a neural network may be partitioned differently by the following layer. When defined in terms of the input space, or $\bm{x}$} space, the hyperplanes associated with the second layer differ across the linear regions defined by the first layer because each of those linear regions is associated with a different affine transformation from $\bm{x}$} to ${\bm{h}}^{1}$. Hence, the activation hyperplanes of layer $l$ may break each linear region from layer l1l-1 differently. To every linear region defined by the hyperplane arrangement in the ${\bm{h}}^{l-1}$ space there is a linear transformation ${\bm{h}}^{l-1}=\Omega^{{\mathbb{S}}^{l-1}}({\bm{W}}^{l-1}{\bm{h}}^{l-2}+{\bm{b}}^{l-1})$ to the points of that linear region based on the set of active neurons ${\mathbb{S}}^{l-1}$. Consequently, inputs in the ${\bm{h}}^{l-1}$ space that are associated with different linear regions are transformed differently to the ${\bm{h}}^{l}$ space, and therefore the form in which those linear regions are further partitioned by layer $l$ is not the same when seen from the ${\bm{h}}^{l-1}$ space.
Second, some combinations of activation sets of multiple layers do not correspond to linear regions even if the activation hyperplanes are in general position with respect to each layer. For each layer, the first two plots of Figure 6 show that every activation set corresponds to a nonempty region of the layer input. However, not every pair of such activation sets would define a nonempty linear region for the neural network. For example, the linear region of the first layer associated with the activation set ${\mathbb{S}}^{1}=\{\}$ defines a linear region in $\bm{x}$} which is always mapped to ${\bm{h}}^{1}=0$, and thus only corresponds to activation set ${\mathbb{S}}^{2}=\{\gamma,\delta\}$ in the second layer because both units are active for such input. Thus, no linear region in $\bm{x}$} is associated with only the units in sets $\{\},\{\gamma\}$, and $\{\delta\}$ being active —i.e., there is no linear region such that ${\mathbb{S}}^{1}\cup{\mathbb{S}}^{2}=\{\},\{\gamma\},\text{or}\{\delta\}$.
More generally, the number of units that are active on each linear region defined by the first layer also imposes a geometric limit to how that linear region can be further partitioned. If only one unit is active at a layer, that means that the output of the layer within that linear region has dimension 1, and, consequently, the subsequent hyperplane arrangements within that linear region are limited to a 1-dimensional space. For the network in the example, we thus expect no more than $\sum_{i=0}^{1}\binom{2}{i}=3$ linear regions being defined instead of $2^{2}=4$ when only one unit is active. In fact, that is precisely the number of subdivisions by the second layer of the linear region defined by activation set ${\mathbb{S}}^{1}=\{\beta\}$ from the first layer.
@@ 3.4 Piecewise linear representability
A natural follow-up question is whether one can do better than this logarithmic bound. Hertrich et al. (2023) conjectured that this is not true.
@@
To narrow down the path into proving this conjecture, Hertrich et al. (2023) show that Conjecture 1 is equivalent to the following.
@@
During the review process of this survey, Conjectures 1 and 2 were disproved by Bakaev et al. (2025).
Theorem 2 (Bakaev et al. (2025))
Every piecewise linear function $\mathbb{R}^{n}\to\mathbb{R}$ can be represented by a ReLU DNN with at most $\lceil\log_{3}(n-1)\rceil+1$ hidden layers.
Other relevant recent works around these conjectures are Valerdi (2024), Averkov et al. (2025), Grillo et al. (2025) and Bakaev et al. (2024). In what follows, we discuss prior work towards these conjectures, which produced notable results. For example, Hertrich et al. (2023) proved the following.
@@
Later on, the conjecture was shown to be true for the case of networks with integer weights.
@@
Consider $f$ a positively homogeneous, convex, piecewise linear function (such as the maximum function). Such a function can be written as $f(x)=\max\{a_{1}^{\top}x,\ldots,a_{p}^{\top}x\}$ for some family of vectors $\{a_{i}\}_{i=1}^{p}$; the Newton polytope $P_{f}$ is defined as $P_{f}:=\mathrm{conv}\{a_{1},\ldots,a_{p}\}$. Note that if $f(x)=\max\{0,x_{1},\ldots,x_{n}\}$ its Newton polytope is the unit simplex, i.e., $P_{f}=\Delta_{0}^{n}:=\mathrm{conv}\{0,e_{1},\ldots,e_{n}\}$ where $e_{i}$ is the $n$-th dimensional $i$-th unit vector.
Haase et al. (2023) leverage a strong connection between Newton polytopes and ReLU networks: that the ++ and $\max$ operations over functions (which are the building blocks in ReLU networks) translate well to operations over polytopes, namely, $P_{f+g}=P_{f}+P_{g}$ and $P_{\max\{f,g\}}=\mathrm{conv}(P_{f}\cup P_{g})$. Then, to connect the functions representable by networks with $k$ hidden layers and families of polytopes, the authors define the following recursive family.
$\mathcal{P}^{\prime}_{k+1}$ $:=\{\mathrm{conv}(Q_{1}\cup Q_{2})\,\mid\,Q_{1},Q_{2}\in\mathcal{P}_{k}\}$
$\mathcal{P}_{k+1}$ $:=\{Q_{1}+\cdots+Q_{\ell}\,\mid\,Q_{1},\ldots,Q_{\ell}\in\mathcal{P}^{\prime}_{k+1}\}.$
@@
A positively homogeneous continuous piecewise linear function $f$ can be represented by an integral $k$ hidden-layer neural network if and only if f=ghf=g-h for two convex, positively homogeneous continuous piecewise linear functions $g$ and $h$ with $P_{g},P_{h}\in\mathcal{P}_{k}$.